OpenStudy (anonymous):

What is the similarity ratio of two cubes with volumes of 125 cubic meters and 216 cubic meters? 125:216 25:36 1:3 5:6

OpenStudy (cwrw238):

take cube root of each number

OpenStudy (anonymous):

how

OpenStudy (cwrw238):

calculator will do it on mine i enter 125 ^ (1/3) =

OpenStudy (anonymous):

i dont get it

OpenStudy (cwrw238):

2*2*2 = 8 cube root of 8 = 2

OpenStudy (phi):

what number times itself 3 times is 125?

OpenStudy (phi):

look on your calculator for\[\sqrt[3]{ }\]

OpenStudy (sasogeek):

a cube is a box with all the sides equal. the volume of a cube is one side times itself 3 times.  volume = x * x * x since there's is two of them, except for each of them, the value of x is different (so for one lets say one side is x, and for the other box the side is y, when u have to find the similarity ratio, u need only one side. to do that u need to find the cube root of the given values. answer = cube root of 125 : cube root of 216 = x : y